### 07 Apr 2024 v1

@@ Abstract
This paper introduces StockGPT, an autoregressive “number” model pretrained directly on the history of daily U.S. stock returns. Treating each return series as a sequence of tokens, the model excels at understanding and predicting the highly intricate stock return dynamics. Instead of relying on handcrafted trading patterns using historical stock prices, StockGPT automatically learns the hidden representations predictive of future returns via its attention mechanism. On a held-out test sample from 2001 to 2023, a daily rebalanced long-short portfolio formed from StockGPT predictions earns an annual return of 119% with a Sharpe ratio of 6.5. The StockGPT-based portfolio completely explains away momentum and long-/short-term reversals, eliminating the need for manually crafted price-based strategies and also encompasses most leading stock market factors. This highlights the immense promise of generative AI in surpassing human in making complex financial investment decisions and illustrates the efficacy of the attention mechanism of large language models when applied to a completely different domain.
@@ 1 Introduction
Since StockGPT makes its return forecasts using historical price data, I examine how it relates to common price-based strategies such as momentum and long-/short-term reversals. I find that the StockGPT-based portfolio completely explains away these strategies via the spanning test. The StockGPT-based portfolio also spans most factors of the [33, ] five factor model and the [37, ] q-factor model. That a strategy based solely on historical price data delivers such a strong future performance poses a strong challenge to the market efficiency hypothesis of [32, ] who argues that the stock market is efficient so historical prices cannot be used to consistently predict future returns.
@@ 2.2 Details
At the heart of the transformer model is the attention mechanism. Accordingly, for each token, the transformer generates three vectors from its embedding vector: key $k$, query $q$, and value $v$. The attention for each token is the weighted sum of its $v$ with all $v$’s of tokens preceding it, weighted by the product of its $q$ with $k$’s of those tokens. Intuitively, a token emits a query and the previous tokens that can match its query (i.e., having a high $q\times k$ value) get its attention. $k$, $q$, and $v$ are also learnable parameters.777 This is a loose use of language since technically speaking, the model learns the weight matrices that produce these vectors. This mechanism constitutes a “self-attention” head and helps the transformer develop a contextual understanding of the tokens. Since each token is only influenced by the tokens before it, this setup is autoregressive.
@@ 2.3 StockGPT Specifics
Besides the vocabulary size of 402, StockGPT has a block size (i.e., length of each input sequence) of 256, token and positional embedding sizes of 128, 4 attention blocks each consisting of 4 self-attention heads, and a dropout probability of 0.2 in various layers. Taken together, StockGPT has 927,890 parameters. StockGPT is trained in 10,000 training steps with each step consisting of 64 sequences (i.e, batch size) drawn randomly from the training data.999 During the training phase, the cross-entropy loss stabilizes at around 2.5 after 5,000 training steps. With 402 labels (the number of return bins), the maximum cross-entropy would be $E=-\sum_{i}log(1/402)\times(1/402)=6$. The model is fully trained locally on a MacBook M2 with 64GB RAM and 30 GPU cores. The probability of sampling each stock during training is proportional to the number of daily return observations it has.
@@ 4.3 Monthly Returns
The second approach is to train a new StockGPT model where the unit of each return observation is $m$ days (e.g., training a StockGPT model using monthly returns as input). I pursue this approach in this section. Specifically, I train a StockGPT model on return sequences of 128 months (approximately 10 years) with all other specifications kept unchanged from the daily model. Since the model now uses monthly returns and each stock has to have at least 128 monthly observations, the training sample is left with only 1.8 million monthly observations. For this model, I include data from all three exchanges (NYSE, NASDAQ, and AMEX) since the 128-month requirement already eliminates most noises from smaller stock exchanges like NASDAQ and AMEX.
The monthly long-short portfolios after removing the bottom 10th percentile stocks based on market value are reported in Table 5. The equal-weighted long-short monthly portfolio earns about 7-8% annually, significant at 1%, with a Sharpe ratio around 0.6-0.7. These equal-weighted portfolios also do not decay quickly if one month is skipped between return forecasts and portfolio formation. Whether one month is skipped or not, the monthly StockGPT models is better at forecasting stocks in the long leg than stocks in the short leg as indicated in Figure 3.
Overall, while the monthly StockGPT model still produces profitable trading strategies, the magnitude of monetary profits is much smaller than that of the daily model. This result is expected for two reasons. First, the amount of training data is much smaller under the monthly approach. Second, the requirement of having at least 10 years of data eliminates most small cap stocks where the super-performance of the daily strategy stems from.
@@ 5 Conclusion
This figure plots the cumulative returns of long-short high-minus-low (HML), high (H), and low (L) portfolios formed from StockGPT return forecasts. Portfolios are equal-weighted. Portfolios are formed after excluding stocks in the bottom decile based on market value. Different portfolios correspond to different market price thresholds under which stocks are excluded. Panel A (B) shows results when return forecasts for month $t+1$ are used to formed portfolios for month $t+1$ ($t+2$). The left (right) panel shows results for the long-short (each leg) portfolios. The sample is monthly from January 2001 to December 2023.
@@
Table 2: Fama-MacBeth Regression
@@
Table 4: Spanning Tests
@@
This table reports the return statistics of the monthly long-short StockGPT-based portfolios. Mean and SD (standard deviation) are in annualized percentage points; Mean/SD (Sharpe ratio) is annualized; Min, Max, and MDD (max drawdown) are in percentage points; and t-Mean is $t$-statistic of the mean portfolio return using Newey-West standard error with 4 lags. Portfolios are formed after excluding stocks in the bottom decile based on market value. Horizon 1 (2) refers to using return forecasts for month $t+1$ to form portfolios for month $t+1$ ($t+2$). EW (VW) refers to equal weighting (value weighting). Price Filter refers to the price level under which stocks are removed. The sample is monthly from January 2001 to December 2023.
Horizon Weight Price Filter Mean t-Mean SD Mean/SD Min Max MDD 1 EW 0 7.9 2.9 13.2 0.6 -11.0 27.8 -22.0 1 EW 1 7.3 3.0 12.1 0.6 -10.8 21.6 -19.2 1 EW 3 7.1 3.4 10.3 0.7 -8.6 13.8 -17.7 1 EW 5 6.3 3.0 10.0 0.6 -8.9 13.5 -16.7 1 VW 0 6.3 1.4 20.8 0.3 -26.3 33.4 -52.9 1 VW 1 5.6 1.4 19.3 0.3 -26.7 22.4 -48.0 1 VW 3 7.0 1.9 18.3 0.4 -27.1 16.4 -39.0 1 VW 5 6.8 1.9 17.5 0.4 -27.4 17.1 -40.0 2 EW 0 6.4 2.5 12.2 0.5 -7.8 29.2 -26.7 2 EW 1 6.8 3.1 10.9 0.6 -7.1 25.4 -21.2 2 EW 3 7.3 3.6 10.3 0.7 -7.0 25.2 -15.3 2 EW 5 6.8 3.3 9.5 0.7 -6.9 20.1 -18.0 2 VW 0 -1.5 -0.3 21.6 -0.1 -29.4 22.3 -75.7 2 VW 1 -1.8 -0.4 21.5 -0.1 -28.6 25.7 -75.0 2 VW 3 -0.2 -0.0 19.0 -0.0 -25.0 22.8 -64.7 2 VW 5 -1.2 -0.3 17.8 -0.1 -21.9 22.2 -66.3

### 09 Apr 2024 v2

@@ Abstract
This paper introduces StockGPT, an autoregressive “number” model trained and tested on 70 million daily U.S. stock returns over nearly 100 years. Treating each return series as a sequence of tokens, StockGPT automatically learns the hidden patterns predictive of future returns via its attention mechanism. On a held-out test sample from 2001 to 2023, a daily rebalanced long-short portfolio formed from StockGPT predictions earns an annual return of 119% with a Sharpe ratio of 6.5. The StockGPT-based portfolio completely spans momentum and long-/short-term reversals, eliminating the need for manually crafted price-based strategies, and also encompasses most leading stock market factors. This highlights the immense promise of generative AI in surpassing human in making complex financial investment decisions.
@@ 1 Introduction
Since StockGPT makes its return forecasts using historical price data, I examine how it relates to common price-based strategies such as momentum and long-/short-term reversals. I find that the StockGPT-based portfolio completely spans these strategies via the spanning test. The StockGPT-based portfolio also spans most factors of the [33, ] five factor model and the [37, ] q-factor model. That a strategy based solely on historical price data delivers such a strong future performance poses a strong challenge to the market efficiency hypothesis of [32, ] who argues that the stock market is efficient so historical prices cannot be used to consistently predict future returns.
@@ 2.2 Details
At the heart of the transformer model is the attention mechanism. Accordingly, for each token, the transformer generates three vectors from its embedding vector: key $k$, query $q$, and value $v$. The attention for each token is the weighted sum of its $v$ with all $v$’s of tokens preceding it, weighted by the product of its $q$ with $k$’s of those tokens. Intuitively, a token emits a query and the previous tokens that can match its query (i.e., having a high $q\times k$ value) get its attention. $k$, $q$, and $v$ are also learnable parameters.777 Technically speaking, the model learns the weight matrices that produce these vectors. This mechanism constitutes a “self-attention” head and helps the transformer develop a contextual understanding of the tokens. Since each token is only influenced by the tokens before it, this setup is autoregressive.
@@ 2.3 StockGPT Specifics
Besides the vocabulary size of 402, StockGPT has a block size (i.e., length of each input sequence) of 256, token and positional embedding sizes of 128, 4 attention blocks each consisting of 4 self-attention heads, and a dropout probability of 0.2 in various layers. Taken together, StockGPT has 0.93 million parameters. StockGPT is trained in 10,000 training steps with each step consisting of 64 sequences (i.e, batch size) drawn randomly from the training data.999 During the training phase, the cross-entropy loss stabilizes at around 2.5 after 5,000 training steps. With 402 labels (the number of return bins), the maximum cross-entropy would be $E=-\sum_{i}log(1/402)\times(1/402)=6$. The model is fully trained locally on a MacBook M2 with 64GB RAM and 30 GPU cores. The probability of sampling each stock during training is proportional to the number of daily return observations it has.
@@ 4.3 Monthly Returns
The second approach is to train a new StockGPT model where the training target is to predict return over the next $m$ days (i.e., $p(\bar{x}_{t+1\rightarrow t+m}|x_{t},\dots,x_{t-255})$ where $\bar{x}_{t+1\rightarrow t+m}$ is mean return over the next $m$ days). I pursue this approach in this section. Specifically, I train a StockGPT model using historical returns to predict mean returns over the next 20 days (i.e., 20 days approximates the number of trading days in a month) with all other specifications kept unchanged from the daily model. During the testing phase, at the end of each month for each stock, an input sequence of 256 previous daily returns for that stock is used by the new StockGPT model to predict the next 20-day returns (i.e., return over the next month).
To evaluate the quality of long-term forecasts by StockGPT, I regress actual realized returns onto 20-day return forecasts via the Fama-McBeth test discussed above. As reported in Table 5, the average slope coefficient is 3 (significant at 5%), indicating that a cross-sectional difference of 100 basis points (i.e, 1%) in StockGPT return forecasts signals a difference of 300 basis points in realized returns. Moreover, the average cross-sectional $R^{2}$ is 0.55% equivalent to a 7.4% cross-sectional correlation between return predictions and actual returns. When one month is skipped between 20-day return forecasts and realized returns, the correlation shrinks to zero.
I then form monthly long-short decile portfolios using the 20-day return forecasts. The monthly long-short portfolios after removing the bottom 10th percentile stocks based on market value are reported in Table 6. The equal-weighted long-short monthly portfolios earn about 13% annually, significant at 1%, with Sharpe ratios around 1. Meanwhile, the value-weighted long-short portfolios earn about 9% to 11% annually, with Sharpe ratios around 0.4-0.5. If one month is skipped between return forecasts and portfolio formation, the returns and Sharpe ratios of the equal-weighted portfolios are halved while those of the value-weighted portfolios remain more stable.
Figure 3 plots the log cumulative returns on the equal-weighted portfolios. The long-short portfolios see a stable upward trend from 2001 to 2023. Consistent with the results in Table 5, when one month is skipped, the portfolios lose their performance, mostly going flat in the second half of the test sample. Whether one month is skipped or not, StockGPT does better at predicting stocks in the long leg.
Overall, while the monthly StockGPT model still produces profitable trading strategies, the magnitude of monetary profits is much smaller than that of the daily model. This suggests that autoregressive models such as StockGPT do the best when predicting the next token in the sequence. Moreover, under monthly rebalancing, we have to wait for one month before updating the portfolios, potentially missing out short-term profitable opportunities.
@@ 5 Conclusion
This figure plots the log cumulative returns of long-short high-minus-low (HML), high (H), and low (L) portfolios formed from StockGPT return forecasts. Portfolios are equal-weighted. Portfolios are formed after excluding stocks in the bottom decile based on market value. Different portfolios correspond to different market price thresholds under which stocks are excluded. Panel A (B) shows results when return forecasts for month $t+1$ are used to formed portfolios for month $t+1$ ($t+2$). The left (right) panel shows results for the long-short (each leg) portfolios. The sample is monthly from February 2001 to December 2023.
@@
Table 2: Daily Fama-MacBeth Regression
@@
Table 4: Daily Spanning Tests
@@
Table 5: Monthly Fama-MacBeth Regression
where $x_{it+1}$ is the actual realized return of stock $i$ in month $t+1$ and $\hat{x}_{it+1}$ is its StockGPT return forecast. Returns are in basis points and $R^{2}$ in percentage points. $t$ is the $t$-statistic of time-series mean of $b_{t}$ computed using [46, ] standard error with 4 lags. Horizon 1 (2) means comparing return forecasts for month $t+1$ with actual returns in month $t+1$ ($t+2$). The sample is monthly from February 2001 to December 2023.
    Horizon         b         $t$         $R^{2}$
    1         3.01         2.49         0.55
    2         -0.08         -0.08         0.43
@@
This table reports the return statistics of the monthly long-short StockGPT-based portfolios. Mean and SD (standard deviation) are in annualized percentage points; Mean/SD (Sharpe ratio) is annualized; Min, Max, and MDD (max drawdown) are in percentage points; and t-Mean is $t$-statistic of the mean portfolio return using Newey-West standard error with 4 lags. Portfolios are formed after excluding stocks in the bottom decile based on market value. Horizon 1 (2) refers to using return forecasts for month $t+1$ to form portfolios for month $t+1$ ($t+2$). EW (VW) refers to equal weighting (value weighting). Price Filter refers to the price level under which stocks are removed. The sample is monthly from February 2001 to December 2023.
Horizon Weight Price Filter Mean t-Mean SD Mean/SD Min Max MDD 1 EW 0 13.5 3.8 14.9 0.9 -14.3 18.4 -35.0 1 EW 1 13.5 4.2 13.5 1.0 -14.3 17.1 -27.2 1 EW 3 13.1 4.1 13.6 1.0 -15.6 12.0 -20.6 1 EW 5 12.5 3.9 13.8 0.9 -14.5 12.7 -25.5 1 VW 0 11.6 2.3 24.4 0.5 -24.8 27.0 -54.7 1 VW 1 8.6 1.7 23.9 0.4 -20.8 26.9 -50.4 1 VW 3 9.6 2.1 22.4 0.4 -20.5 20.2 -39.5 1 VW 5 9.1 2.0 21.5 0.4 -20.7 18.1 -37.2 2 EW 0 7.5 2.0 15.8 0.5 -15.3 19.0 -27.9 2 EW 1 6.3 2.0 14.0 0.4 -18.3 15.7 -28.3 2 EW 3 5.7 1.8 13.8 0.4 -17.1 11.0 -32.9 2 EW 5 5.6 1.8 13.9 0.4 -17.2 12.7 -35.4 2 VW 0 12.6 2.1 28.2 0.4 -36.7 32.1 -61.8 2 VW 1 9.8 1.8 26.9 0.4 -36.7 25.6 -62.7 2 VW 3 7.3 1.5 24.5 0.3 -33.5 19.8 -64.2 2 VW 5 5.7 1.2 23.5 0.2 -30.4 19.4 -61.7